# Softpick: No Attention Sink, No Massive Activations with Rectified Softmax

### v1

@@ 6.2 Attention Sink
Attention sink can be identified by the large scores on the first column in the attention map, which manifests as a bright vertical line leftmost on the attention heatmaps. Notably, the heads that have heavy attention sinks i.e. large scores on the BOS token and close to none on other tokens with softmax, become either specifically tuned to only score specific trigger tokens, or are completely shut off when not needed. An example of this can be seen in Figure 5, on head 4 of the 10th layer. The softmax model has a clear attention sink and barely attends to other tokens, while the softpick model completely shuts off the head (all zero scores) on the first input, and attends to some specific tokens on the second and third input sample text. We believe this behavior is a more concrete version of the active-dormant heads in vanilla attention, a previously observed phenomenon [Guo et al., 2024]. We discuss this and its implications in 7.4.
In addition to visually observing attention sinks, we can use the sink rate metric proposed in [Gu et al., 2024], defined as $\frac{1}{L}\sum_{l=1}^{L}\frac{1}{H}\sum_{h=1}^{H}\mathbb{I}(\alpha_{1}^{l,h}>\epsilon_{s})$ which is the percentage of heads in the transformer that on average has an attention score above some threshold $\epsilon_{s}$ in the first token, where $\alpha_{1}^{l,h}$ is the average value of the first column in $\mathbf{A}$ at layer $l$ and head $h$. We calculate the sink rate of both models on 1000 samples from the validation set of SlimPajama [Soboleva et al., 2023] with $\epsilon_{s}=0.3$ (as suggested in [Gu et al., 2024]) and also a stricter $\epsilon_{s}=0.6$ and present the results in Table 3. The model using softpick has a sink rate of 0% for both thresholds, which means that no attention heads are overscoring the BOS token in any scenario, effectively eliminating attention sink completely. The effects of this can be seen when we analyze the hidden activations between transformer layers and the large values inside them.
@@ 6.3 Massive Activations
We analyze the effect of softpick on massive activations by examining the hidden state outputs after every transformer layer. We present some key metrics in Table 3 and box plots of the hidden state activation distribution of every two layers (+ the last layer) in Figure 4. These metrics are calculated by running 10 samples of text through the model and collecting all hidden state vectors to evaluate them all at once. The box plots visualize quite clearly how large the massive activations of the softmax model are, so much so as to making the boxes (Q1, median, and Q3) barely visible due to the large minimum and maximum values. This is especially prominent from the middle layers up to second to last layer. Meanwhile, the softpick models do not exhibit these massive activations. There is an exception in the last layer where it matches the softmax model, most likely due to the need to project back into the vocabulary space. This difference in activations is quite dramatic and can be seen clearly in the metrics as well. Kurtosis of all hidden state activations is significantly reduced from 33,510 to 340, a hundred-fold reduction. The minimum and maximum values are also reduced by an order of magnitude. Note that the minimum and maximum values of the softpick model are in fact from the last layer, all previous layers have even smaller activations.
@@ 7.1 Quantization
Quantizing models presents challenges due to a scaling problem caused by large values within the activations [Dettmers et al., 2022]. Naive quantization of these values leads to inaccurate scaling since simply quantizing a vector containing such values leads to bias towards the maximum scores from the massive activation, which makes non-outliers underrepresented. A common solution is to use blockwise absmax quantization [Dettmers et al., 2022, 2023]; however, this method requires storing numerous maximum values. QLoRA [Dettmers et al., 2023] mitigates this storage requirement through double quantization, which reduces the storage requirement for the maximum states for each block, but this introduces approximations, as the calculations are not exact, potentially reducing accuracy. We propose that removing attention sinks, identified as significant outliers in activation, should help address this scaling problem. As shown in Figure 3, our method, softpick, outperforms softmax across various benchmarks and quantization techniques. The performance advantage of softpick becomes particularly pronounced when lower bit precision (lower qbit) is employed. Softpick also opens up the avenue for new quantization algorithms in the future that do not need to accommodate massive activations.
@@ 7.2 Low-Precision Training
Low-precision training [Wang et al., 2025, Fishman et al., 2024, Peng et al., 2023, DeepSeek-AI et al., 2025] is an emerging topic, driven by recent hardware supporting FP8 [NVIDIA, b] and FP4 [NVIDIA, a] Tensor Cores, which offer higher FLOPS. Peng et al. [2023] found that lower precision can be used for gradients and optimizers without significant accuracy loss or changes to hyperparameters. Fishman et al. [2024] identified that sudden outliers in activations, potentially amplified by functions such as SwiGLU [Shazeer, 2020], disrupt the statistical assumptions needed for low-bit training and can cause divergence. They addressed this by introducing Smooth-SwiGLU which also enables them to quantize the optimizer’s second moment. To handle outliers and massive activations, Wang et al. [2025] applied online Hadamard transformations to make activations more Gaussian-like, enabling aggressive low-bit training. Similarly, DeepSeek-AI et al. [2025] created a fine-grained quantization scheme to reduce quantization errors caused by large activations.
Since outliers are a primary challenge for low-bit training [Fishman et al., 2024], our method aims to produce activations without large activations or attention sink. We hypothesize that by removing this main obstacle, our approach will potentially allow simple low-bit training schemes to be used directly without issue, mitigating the divergence problem that occurs when massive activations emerge during training.
@@ 7.3 Sparsity
As discussed in section 6.1, softpick enables sparse attention. This sparsity offers potential for acceleration of inference through sparse GEMM kernels [Qin et al., 2020]. Furthermore, this sparsity allows computational savings by skipping operations involving zero-valued scores, conceptually similar to optimizations based on activation sparsity [Liu et al., 2025]. Specifically in our method, because the sparsity is guaranteed in the attention map (post-softpick), the matrix multiplication with the value matrix $\mathbf{V}$ only needs to be computed for the non-zero attention scores, potentially eliminating redundant calculations.
@@ 7.4 Active-Dormant Heads and Pruning
As outlined in Section 6.1, our observations indicate that some heads in the softpick models stay entirely inactive until activated by particular trigger tokens. As noted in [Guo et al., 2024], these heads are called active-dormant heads due to their behavior. In vanilla softmax transformers, these heads are identified by a large attention sink, a "drained" value state where the value projections are mostly zero, and a large residual activation. We believe that the active-dormant heads in softpick fill a similar purpose, but without the drawbacks that we see in softmax models. This opens up some avenues in head pruning [Wang et al., 2021, Shim et al., 2021], as it is now much easier to identify dormant heads and prune them, without affecting the computation at all if that head would be outputting zero scores anyways.
@@ 7.5 Interpretability
Our attention map analysis shows that the attention maps produced by softpick are much more legible compared to those produced by softmax. Objectively, this is reflected in the sparsity metric discussed previously, but it can also be argued that the attention maps shown in Appendix D are visually more interpretable for softpick. This may have implications for interpretability techniques such as attention rollout [Abnar and Zuidema, 2020].
@@ 7.6 Other Modalities
Attention sink and massive activation outliers have also been observed in visual transformers. Darcet et al. [2024] identified artifacts within the feature maps of the Vision Transformers (ViT) [Dosovitskiy et al., 2021] architecture. They addressed this problem by introducing register tokens designed to function as dedicated sinks. This is similar to adding attention bias in language models. Using softpick in ViTs might alleviate this issue and remove the need to use trainable register tokens that require extra parameters and computation. Attention sink has also been observed in video diffusion transformers [Wen et al., 2025] and streaming ASR models [Kumar et al., 2024] where our function might help. Softpick could also potentially be useful in removing attention sinks in multimodal models.
@@ 7.7 Long Context and The Issue of Underscoring
Table 4: Passkey retrieval results in terms of accuracy at certain sequence lengths.
@@
One of the potential benefits of sharper and sparse attention scores is the reduction of noise when accumulating values. Unlike softmax where every value must have a non-zero score, softpick can assign zero to unneeded values, theoretically leading to better retrieval for long context use. However, we have yet to see these benefits in our testing. We show some results on the passkey retrieval task [Lu et al., 2024] of both models on several sequence lengths in Table 4. Overall, softpick does not outperform softmax, at least not on this task.
We have a hypothesis as to why this is the case. The main issue is that softpick assigns smaller scores the more sparse the attention pattern and the longer the context is. Due to the nature of the denominator containing the absolute value function, if the long context has mostly negative scored tokens, those will significantly scale down the numerator on the few positive scored tokens. This can be seen in several of the attention maps in Appendix D, where the more retrieval-specific heads that only attend to singular tokens have a much smaller scale in scores (see the color bar of the heatmaps). This might result in too weak of a signal from the values, resulting in weak retrieval.
One possible solution that we have considered is to scale up the input scores i.e. the query-key dot product result before going into softpick. This is an approach first proposed in Scalable-Softmax [Nakanishi, 2025] and recently used in Llama 4 models [Meta, 2023]. However, we did not see significant improvement on passkey retrieval results and observed worse training loss and downstream task performance. We detail our experiments and results on this "scalable-softpick" approach in Appendix B. It remains an open problem to solve this issue, and we will continue to look for solutions.
8 Future Additions
Scaling up
We are planning to scale up our experiments on up to 2B and 7B parameter scales in the next versions of this paper, to confirm whether softpick works at scale or not.
More analysis
We will add more analysis, especially on the behavior of attention sinks and massive activations during training, with and without softpick.
Solving the issue of underscoring
We will attempt to solve the underscoring issue in the hope of improving scores on long-context tasks.
@@ Acknowledgments
We would like to thank fal.ai and their grant program (https://fal.ai/grants) for providing us with the GPU hours needed to train all 340M models.
@@ Appendix A FlashAttention Algorithms
9:    On chip, compute $m_{i}^{(j)}=\mathrm{max}(m_{i}^{(j-1)},\mathrm{rowmax}(\mathbf{S}_{i}^{(j)}))\in\mathbb{R}^{B_{r}}$$\tilde{\mathbf{P}}_{i}^{(j)}=\exp(\mathbf{S}_{i}^{(j)}-m_{i}^{(j)})-\exp(-m_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\tilde{\mathbf{R}}_{i}^{(j)}=ReLU(\tilde{\mathbf{P}}_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\tilde{\mathbf{A}}_{i}^{(j)}=|\tilde{\mathbf{P}}_{i}^{(j)}|\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\ell_{i}^{(j)}=e^{m_{i}^{j-1}-m_{i}^{(j)}}\ell_{i}^{(j-1)}+\mathrm{rowsum}(\tilde{\mathbf{A}}_{i}^{(j)})\in\mathbb{R}^{B_{r}}$.
@@ Appendix B Other experiments
$\nu(x)$ $=\begin{cases}x&x\geq 0\\ -\infty&x<0\end{cases}$
@@
Second is the "scalable-softpick" approach, an attempt to improve long context performance taking reference from [Nakanishi, 2025]:
@@
where $s$ is a per-head trainable parameter and $n$ is the sequence length. Note that in the training code, $n$ is a vector which assigns a different sequence length, the corresponding one, for each row i.e. token position. We train both models on the same setup and configuration as our two main models. We report the training loss and gradient norm plots in Figure 6, the benchmark results in Table 5, and passkey retrieval results in Table 6.
@@ Appendix C Full Result on Quantization
Table 7: Comparison of softpick vs softmax Performance for BNB Dettmers et al. [2022, 2023] Quantization Methods. $\uparrow$=Higher is Better, $\downarrow$=Lower is Better. $\Delta$ = SoftPick - Vanilla.
Task Metric Quantization Softmax Softpick $\Delta$ (S - V)
Arc Easy Acc Norm $\uparrow$ 4-bit 53.70 53.96 +0.26
Acc $\uparrow$ 4-bit 59.01 59.22 +0.21
Acc Norm $\uparrow$ 8-bit 56.06 56.61 +0.55
Acc $\uparrow$ 8-bit 60.23 60.77 +0.54
Lambada Acc $\uparrow$ 4-bit 33.59 31.61 -1.98
Perplexity $\downarrow$ 4-bit 37.60 39.26 +1.66
Acc $\uparrow$ 8-bit 35.78 35.96 +0.18
Perplexity $\downarrow$ 8-bit 31.30 28.69 -2.61
Piqa Acc Norm $\uparrow$ 4-bit 65.18 66.32 +1.14
Acc $\uparrow$ 4-bit 65.07 66.10 +1.03
Acc Norm $\uparrow$ 8-bit 66.54 66.87 +0.33
Acc $\uparrow$ 8-bit 67.08 66.27 -0.81
Sciq Acc Norm $\uparrow$ 4-bit 75.00 75.10 +0.10
Acc $\uparrow$ 4-bit 83.10 83.00 -0.10
Acc Norm $\uparrow$ 8-bit 75.40 77.10 +1.70
Acc $\uparrow$ 8-bit 82.70 83.40 +0.70
Wikitext Word Perplexity $\downarrow$ 4-bit 223.39 153.46 -69.93
8-bit 205.51 142.94 -62.57
@@
Table 8: Comparison of softpick vs softmax performance for GPTQ Frantar et al. [2023] quantization. $\uparrow$=Higher is Better, $\downarrow$=Lower is Better. $\Delta$ = SoftPick - Vanilla.
Task Metric Quantization Softmax Softpick $\Delta$ (S - V)
Arc Easy Acc Norm $\uparrow$ 2-bit 27.02 29.59 +2.57
Acc $\uparrow$ 2-bit 26.77 29.67 +2.90
Acc Norm $\uparrow$ 3-bit 48.57 51.73 +3.16
Acc $\uparrow$ 3-bit 53.79 56.57 +2.78
Acc Norm $\uparrow$ 4-bit 53.75 54.17 +0.42
Acc $\uparrow$ 4-bit 60.14 59.47 -0.67
Acc Norm $\uparrow$ 8-bit 56.10 56.57 +0.47
Acc $\uparrow$ 8-bit 60.31 61.03 +0.72
Lambada Acc $\uparrow$ 2-bit 0.04 0.39 +0.35
Perplexity $\downarrow$ 2-bit 925599.68 133234.54 -792365.14
Acc $\uparrow$ 3-bit 23.71 28.97 +5.26
Perplexity $\downarrow$ 3-bit 82.08 52.41 -29.67
Acc $\uparrow$ 4-bit 33.61 35.05 +1.44
Perplexity $\downarrow$ 4-bit 37.29 31.04 -6.24
Acc $\uparrow$ 8-bit 36.52 36.21 -0.31
Perplexity $\downarrow$ 8-bit 30.06 28.66 -1.40
Piqa Acc Norm $\uparrow$ 2-bit 48.86 53.92 +5.06
Acc $\uparrow$ 2-bit 52.01 54.79 +2.78
Acc Norm $\uparrow$ 3-bit 63.71 64.69 +0.98
Acc $\uparrow$ 3-bit 64.74 65.61 +0.87
Acc Norm $\uparrow$ 4-bit 65.89 66.65 +0.76
Acc $\uparrow$ 4-bit 65.83 66.43 +0.60
Acc Norm $\uparrow$ 8-bit 66.76 66.76 +0.00
Acc $\uparrow$ 8-bit 67.25 66.59 -0.66
Sciq Acc Norm $\uparrow$ 2-bit 29.80 36.00 +6.20
Acc $\uparrow$ 2-bit 27.70 35.10 +7.40
Acc Norm $\uparrow$ 3-bit 73.80 75.90 +2.10
Acc $\uparrow$ 3-bit 79.40 82.40 +3.00
Acc Norm $\uparrow$ 4-bit 73.50 76.80 +3.30
Acc $\uparrow$ 4-bit 80.70 83.10 +2.40
Acc Norm $\uparrow$ 8-bit 75.10 77.50 +2.40
Acc $\uparrow$ 8-bit 83.10 83.50 +0.40
Wikitext Word Perplexity $\downarrow$ 2-bit 17118.42 2520.87 -14597.56
3-bit 259.32 181.12 -78.21
4-bit 213.65 148.96 -64.69
8-bit 204.34 141.49 -62.85

### v2

@@ 6.2 Attention Sink
Attention sink can be identified by the large scores on the first column in the attention map, which manifests as a bright vertical line leftmost on the attention heatmaps. Notably, the heads that have heavy attention sinks i.e. large scores on the BOS token and close to none on other tokens with softmax, become either specifically tuned to only score specific trigger tokens, or are completely shut off when not needed. An example of this can be seen in Figure 6, on head 4 of the 10th layer. The softmax model has a clear attention sink and barely attends to other tokens, while the softpick model completely shuts off the head (all zero scores) on the first input, and attends to some specific tokens on the second and third input sample text. We believe this behavior is a more concrete version of the active-dormant heads in vanilla attention, a previously observed phenomenon [20]. We discuss this and its implications in 7.4.
In addition to visually observing attention sinks, we can use the sink rate metric proposed in [19], defined as $\frac{1}{L}\sum_{l=1}^{L}\frac{1}{H}\sum_{h=1}^{H}\mathbb{I}(\alpha_{1}^{l,h}>\epsilon_{s})$ which is the percentage of heads in the transformer that on average has an attention score above some threshold $\epsilon_{s}$ in the first token, where $\alpha_{1}^{l,h}$ is the average value of the first column in $\mathbf{A}$ at layer $l$ and head $h$. We calculate the sink rate of both models on 1000 samples from the validation set of SlimPajama [42] with $\epsilon_{s}=0.3$ (as suggested in [19]) and also $\epsilon_{s}=0.2$ and present the results in Table 4. The model using softpick has a sink rate of 0% for both thresholds and both model sizes, which means that no attention heads are overscoring the BOS token in any scenario, effectively eliminating attention sink completely. The effects of this can be seen when we analyze the hidden activations between transformer layers and the large values inside them.
@@ 6.3 Massive Activations
We analyze the effect of softpick on massive activations by examining the hidden state outputs after every transformer layer. We present some key metrics in Table 4 and box plots of the hidden state activation distribution of every some layers of the 340M and 1.8B models in Figure 5. These metrics are calculated by running 10 samples of text through the model and collecting all hidden state vectors to evaluate them all at once. The box plots visualize quite clearly how large the massive activations of the softmax model are, so much so as to making the boxes (Q1, median, and Q3) barely visible due to the large minimum and maximum values. This is especially prominent from the middle layers up to second to last layer. Meanwhile, the softpick models do not exhibit these massive activations. There is an exception in the last layer where it matches the softmax model, most likely due to the need to project back into the vocabulary space. This difference in activations is quite dramatic and can be seen clearly in the metrics as well. Kurtosis of all hidden state activations in the 340M model is significantly reduced from 33,510 to 340, a hundred-fold reduction. The minimum and maximum values are also reduced by an order of magnitude. Note that the minimum and maximum values of the softpick model are in fact from the last layer, all previous layers have even smaller activations.
@@ 7.1 Quantization
Quantizing models presents challenges due to a scaling problem caused by large values within the activations [13]. Naive quantization of these values leads to inaccurate scaling since simply quantizing a vector containing such values leads to bias towards the maximum scores from the massive activation, which makes non-outliers underrepresented. A common solution is to use blockwise absmax quantization [13, 14]; however, this method requires storing numerous maximum values. QLoRA [14] mitigates this storage requirement through double quantization, which reduces the storage requirement for the maximum states for each block, but this introduces approximations, as the calculations are not exact, potentially reducing accuracy. It has been proposed that removing attention sinks, identified as significant outliers in activation, should help address this scaling problem. As shown in Figure 3, our 340M softpick model, outperforms softmax across various benchmarks and quantization techniques. The performance advantage of softpick becomes particularly pronounced when lower bit precision (lower qbit) is employed. Softpick also opens up the avenue for new quantization algorithms in the future that do not need to accommodate massive activations.
@@ 7.2 Low-Precision Training
Low-precision training [48, 16, 35, 12] is an emerging topic, driven by recent hardware supporting FP8 [32] and FP4 [31] Tensor Cores, which offer higher FLOPS. Peng et al. [35] found that lower precision can be used for gradients and optimizers without significant accuracy loss or changes to hyperparameters. Fishman et al. [16] identified that sudden outliers in activations, potentially amplified by functions such as SwiGLU [38], disrupt the statistical assumptions needed for low-bit training and can cause divergence. They addressed this by introducing Smooth-SwiGLU which also enables them to quantize the optimizer’s second moment. To handle outliers and massive activations, Wang et al. [48] applied online Hadamard transformations to make activations more Gaussian-like, enabling aggressive low-bit training. Similarly, DeepSeek-AI et al. [12] created a fine-grained quantization scheme to reduce quantization errors caused by large activations.
Since outliers are a primary challenge for low-bit training [16], our method aims to produce activations without large activations or attention sink. We hypothesize that by removing this main obstacle, our approach will potentially allow simple low-bit training schemes to be used directly without issue, mitigating the divergence problem that occurs when massive activations emerge during training.
@@ 7.3 Sparsity
As discussed in section 6.1, softpick enables sparse attention. This sparsity offers potential for acceleration of inference through sparse GEMM kernels [36]. Furthermore, this sparsity allows computational savings by skipping operations involving zero-valued scores, conceptually similar to optimizations based on activation sparsity [24]. Specifically in our method, because the sparsity is guaranteed in the attention map (post-softpick), the matrix multiplication with the value matrix $\mathbf{V}$ only needs to be computed for the non-zero attention scores, potentially eliminating redundant calculations.
@@ 7.4 Active-Dormant Heads and Pruning
As outlined in Section 6.1, our observations indicate that some heads in the softpick models stay entirely inactive until activated by particular trigger tokens. As noted in [20], these heads are called active-dormant heads due to their behavior. In vanilla softmax transformers, these heads are identified by a large attention sink, a "drained" value state where the value projections are mostly zero, and a large residual activation. We believe that the active-dormant heads in softpick fill a similar purpose, but without the drawbacks that we see in softmax models. This opens up some avenues in head pruning [47, 40], as it is now much easier to identify dormant heads and prune them, without affecting the computation at all if that head would be outputting zero scores anyways.
@@ 7.5 Interpretability
Our attention map analysis shows that the attention maps produced by softpick are much more legible compared to those produced by softmax. Objectively, this is reflected in the sparsity metric discussed previously, but it can also be argued that the attention maps shown in Appendix E are visually more interpretable for softpick. This may have implications for interpretability techniques such as attention rollout [1].
@@ 7.6 Other Modalities
Attention sink and massive activation outliers have also been observed in visual transformers. Darcet et al. [11] identified artifacts within the feature maps of the Vision Transformers (ViT) [15] architecture. They addressed this problem by introducing register tokens designed to function as dedicated sinks. This is similar to adding attention bias in language models. Using softpick in ViTs might alleviate this issue and remove the need to use trainable register tokens that require extra parameters and computation. Attention sink has also been observed in video diffusion transformers [50] and streaming ASR models [22] where our function might help. Softpick could also potentially be useful in removing attention sinks in multimodal models.
@@ 7.7 Long Context and The Issue of Underscoring
Table 5: Passkey retrieval results of softmax vs softpick 340M models in terms of accuracy at certain sequence lengths.
@@
One of the potential benefits of sharper and sparse attention scores is the reduction of noise when accumulating values. Unlike softmax where every value must have a non-zero score, softpick can assign zero to unneeded values, theoretically leading to better retrieval for long context use. However, we have yet to see these benefits in our testing. We show some results on the passkey retrieval task [26] of both models on several sequence lengths in Table 5. Overall, softpick does not outperform softmax, at least not on this task.
We have a hypothesis as to why this is the case. The main issue is that softpick assigns smaller scores the more sparse the attention pattern and the longer the context is. Due to the nature of the denominator containing the absolute value function, if the long context has mostly negative scored tokens, those will significantly scale down the numerator on the few positive scored tokens. This can be seen in several of the attention maps in Appendix E, where the more retrieval-specific heads that only attend to singular tokens have a much smaller scale in scores (see the color bar of the heatmaps). This might result in too weak of a signal from the values, resulting in weak retrieval. We also hypothesize that this is one of the reasons why the larger 1.8B model with softpick does not perform as well as softmax in general.
One possible solution that we have considered is to scale up the input scores i.e. the query-key dot product result before going into softpick. This is an approach first proposed in Scalable-Softmax [30] and recently used in Llama 4 models [28]. However, we did not see significant improvement on passkey retrieval results and observed worse training loss and downstream task performance. We detail our experiments and results on this "scalable-softpick" approach in Appendix B. It remains an open problem to solve this issue and find a softmax alternative that scales better than softpick.
@@ Acknowledgments
We would like to thank fal.ai and their grant program (https://fal.ai/grants) for providing us with the GPU hours needed to train all 340M and 1.8B models.
@@ Appendix A FlashAttention Algorithms
9:         On chip, compute $m_{i}^{(j)}=\mathrm{max}(m_{i}^{(j-1)},\mathrm{rowmax}(\mathbf{S}_{i}^{(j)}))\in\mathbb{R}^{B_{r}}$,$\tilde{\mathbf{P}}_{i}^{(j)}=\exp(\mathbf{S}_{i}^{(j)}-m_{i}^{(j)})-\exp(-m_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\tilde{\mathbf{R}}_{i}^{(j)}=ReLU(\tilde{\mathbf{P}}_{i}^{(j)})\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\tilde{\mathbf{A}}_{i}^{(j)}=|\tilde{\mathbf{P}}_{i}^{(j)}|\in\mathbb{R}^{B_{r}\times B_{c}}$ (pointwise), $\ell_{i}^{(j)}=e^{m_{i}^{j-1}-m_{i}^{(j)}}\ell_{i}^{(j-1)}+\mathrm{rowsum}(\tilde{\mathbf{A}}_{i}^{(j)})\in\mathbb{R}^{B_{r}}$.
@@ Appendix B Other experiments
$\nu(x)$ =$=\begin{cases}x&x\geq 0\\ -\infty&x<0\end{cases}$
@@
Second is the "scalable-softpick" approach, an attempt to improve long context performance taking reference from [30]:
@@
where $s$ is a per-head trainable parameter and $n$ is the sequence length. Note that in the training code, $n$ is a vector which assigns a different sequence length, the corresponding one, for each row i.e. token position. We train both models on the same setup and configuration as our two main 340M parameter models. We report the training loss and gradient norm plots in Figure 7, the benchmark results in Table 6, and passkey retrieval results in Table 7.
@@ Appendix C Full Result on Quantization
Table 8: Comparison of softpick vs softmax performance for BNB [13, 14] quantization methods. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 4-bit 53.66 53.87 +0.21
Acc $\uparrow$ 4-bit 58.88 59.60 +0.72
Acc Norm $\uparrow$ 8-bit 55.51 56.36 +0.84
Acc $\uparrow$ 8-bit 59.81 60.82 +1.01
Lambada Acc $\uparrow$ 4-bit 33.63 31.73 -1.90
Perplexity $\downarrow$ 4-bit 37.59 39.25 +1.67
Acc $\uparrow$ 8-bit 35.53 35.86 +0.33
Perplexity $\downarrow$ 8-bit 31.22 28.74 -2.48
Piqa Acc Norm $\uparrow$ 4-bit 65.29 66.38 +1.09
Acc $\uparrow$ 4-bit 65.40 66.32 +0.92
Acc Norm $\uparrow$ 8-bit 66.32 66.32 +0.00
Acc $\uparrow$ 8-bit 67.36 66.38 -0.98
Sciq Acc Norm $\uparrow$ 4-bit 74.90 75.10 +0.20
Acc $\uparrow$ 4-bit 82.90 82.80 -0.10
Acc Norm $\uparrow$ 8-bit 75.20 77.40 +2.20
Acc $\uparrow$ 8-bit 83.40 83.50 +0.10
Wikitext Word Perplexity $\downarrow$ 4-bit 26.38 26.47 +0.09
8-bit 23.95 24.40 +0.46
@@
Table 9: Comparison of softpick vs softmax performance for GPTQ [17] quantization. $\uparrow$=Higher is better, $\downarrow$=Lower is better. $\Delta$ = Softpick - Softmax.
Task Metric Quantization Softmax Softpick $\Delta$
Arc Easy Acc Norm $\uparrow$ 2-bit 27.27 29.76 +2.48
Acc $\uparrow$ 2-bit 27.02 29.88 +2.86
Acc Norm $\uparrow$ 3-bit 48.57 51.81 +3.24
Acc $\uparrow$ 3-bit 53.83 56.65 +2.82
Acc Norm $\uparrow$ 4-bit 53.96 54.29 +0.34
Acc $\uparrow$ 4-bit 59.93 59.26 -0.67
Acc Norm $\uparrow$ 8-bit 56.31 56.36 +0.04
Acc $\uparrow$ 8-bit 60.31 60.94 +0.63
Lambada Acc $\uparrow$ 2-bit 0.04 0.43 +0.39
Perplexity $\downarrow$ 2-bit 930980.10 133928.18 -797051.92
Acc $\uparrow$ 3-bit 23.71 29.07 +5.36
Perplexity $\downarrow$ 3-bit 82.12 52.32 -29.80
Acc $\uparrow$ 4-bit 33.55 34.97 +1.42
Perplexity $\downarrow$ 4-bit 37.30 31.08 -6.21
Acc $\uparrow$ 8-bit 36.44 36.15 -0.29
Perplexity $\downarrow$ 8-bit 30.06 28.66 -1.41
Piqa Acc Norm $\uparrow$ 2-bit 48.42 53.59 +5.17
Acc $\uparrow$ 2-bit 52.12 54.52 +2.39
Acc Norm $\uparrow$ 3-bit 63.76 64.64 +0.87
Acc $\uparrow$ 3-bit 64.53 65.56 +1.03
Acc Norm $\uparrow$ 4-bit 66.10 66.43 +0.33
Acc $\uparrow$ 4-bit 65.89 66.59 +0.71
Acc Norm $\uparrow$ 8-bit 66.59 66.54 -0.05
Acc $\uparrow$ 8-bit 66.81 66.21 -0.60
Sciq Acc Norm $\uparrow$ 2-bit 29.80 35.80 +6.00
Acc $\uparrow$ 2-bit 27.80 34.80 +7.00
Acc Norm $\uparrow$ 3-bit 73.50 75.90 +2.40
Acc $\uparrow$ 3-bit 79.30 82.30 +3.00
Acc Norm $\uparrow$ 4-bit 73.40 76.40 +3.00
Acc $\uparrow$ 4-bit 81.00 82.70 +1.70
Acc Norm $\uparrow$ 8-bit 75.00 77.60 +2.60
Acc $\uparrow$ 8-bit 83.20 83.70 +0.50
Wikitext Word Perplexity $\downarrow$ 2-bit 4081.76 893.35 -3188.41
3-bit 34.57 31.53 -3.04
4-bit 25.69 25.53 -0.16
8-bit 23.84 24.31 +0.47